# A uniform-in-time nonlocal approximation of the standard Fokker-Planck equation

### v2

@@ 4 Convergence: from nonlocal to local
By standard theory of Fokker-Planck equation, the solution $v=v(t,x)$ of (FP) is in $\mathcal{C}^{4}$, for every $t\in[0,T^{*}]$ and $|D^{\alpha}v(t,x)|\leq c\left\langle x\right\rangle^{-K}$ uniformly in $[0,T^{*}]$, due to our assumptions on the derivatives of $u_{0}$. Hence, $v$ satisfies the hypotheses of Proposition 4.1, uniformly for all $t\in[0,T^{*}]$. We rewrite
$Lv=L_{\varepsilon}v+(L_{0}v-L_{\varepsilon}v):=L_{\varepsilon}v+h(t).$
@@
Proposition 4.6 (Convergence of the equilibria).
@@
$\left\|S_{t}\right\|\leq Me^{-\gamma t}\qquad\text{ for }\gamma>0,M>1,t>0$
[…]
@@
where $R$ denotes the resolvent operator. Choosing $n=1$ and $\lambda=0>-\gamma$ in (23), from Theorem 1.1 it follows that $L_{\varepsilon}$ has an inverse with $\left\|L_{\varepsilon}^{-1}\right\|\leq M/\gamma$. Then,
$\begin{split}\left\|F_{\varepsilon}-G\right\|_{L^{1}_{k}}&=\left\|L_{\varepsilon}^{-1}L_{\varepsilon}(F_{e}-G)\right\|\leq\left\|L_{\varepsilon}^{-1}\right\|\cdot\left\|L_{\varepsilon}F_{\varepsilon}-L_{\varepsilon}G\right\|_{L^{1}_{k}}\\ &\leq\frac{M}{\gamma}\left\|L_{\varepsilon}G\right\|_{L^{1}_{k}}=\frac{M}{\gamma}\left\|(L_{\varepsilon}-L_{0})G\right\|_{L^{1}_{k}}.\end{split}$
Given the regularity and decay properties of the Gaussian, by consistency 4.1 we obtain
@@ 5.1 Regularity
Thanks to the the explicit form (11) of the equilibrium in the frequency domain, one can study the regularity by standard results on the Fourier transform.
@@
Recall that the Fourier transform of the equilibrium is
with $\zeta_{\varepsilon}(\xi)=\hat{J}_{\varepsilon}(\xi)-1$. Let us assume $\zeta_{\varepsilon}$ radially symmetric. If not, define $\underline{\zeta}_{\varepsilon}$ and $\overline{\zeta}_{\varepsilon}$ by
@@
Since $\zeta_{\varepsilon}(y)=-\varepsilon^{2}y^{2}+o(y^{2})$, the singularity at zero is cancelled and, moreover, close to $\xi=0$ we have
@@
    which implies that
    $|\hat{F}_{\varepsilon}(\xi)|\leq C\exp\Big{(}-\frac{1-\delta}{\varepsilon^{2}}\log|\xi|\Big{)}=\frac{C}{|\xi|^{\frac{1-\delta}{\varepsilon^{2}}}}\leq\frac{C}{|\xi|^{k+(\frac{1}{\varepsilon^{2}}-k)/2}}.$
    For $|\xi|$ large, the previous implies condition (24) with $\eta<(\frac{1}{\varepsilon^{2}}-k)/2$ and by standard argument on Fourier transform $F_{\varepsilon}\in\mathcal{C}^{k-d}$ for all $k<\frac{1}{\varepsilon^{2}}$.
@@
    proceeding in a similar way as before. Notice that since $\xi\mapsto\hat{J}(\xi)$ is continuous and decays to zero as $|\xi|\to+\infty$, there exists $y_{\delta}$ such that we can bound $\hat{J}$ from below by
    $\hat{J}(x)\geq-\delta:=\frac{\varepsilon^{2}}{2}\Big{(}\frac{1}{\varepsilon^{2}}-k\Big{)}\qquad\text{ for all }y\geq y_{\delta}.$
    As a consequence we see that
@@
Remark 5.2.
We just pointed out that the regularity keeps improving when $\varepsilon\to 0$ although the result is not optimal. It would be interesting to study the regularity of the equilibrium which is, as far as we know, an open question.
Hence we have proved that the regularity of the equilibrium keeps improving as $\varepsilon\to 0$. In fact, we suspect that under quite general conditions, the only point at which $F_{\varepsilon}$ is not $\mathcal{C}^{\infty}$ is the origin. We are able to show this in dimension $1$, and for radially symmetric $J$:
@@
    If $d\neq 1$, the fact that $\operatorname{div}(xF_{\varepsilon})$ is integrable in $\Omega$ does not assure that $\nabla F\in L^{1}(\Omega)$, even when we are far from 00. Therefore, we need an additional hypothesis. If $J$ is radially symmetric, $F_{\varepsilon}$ inherits this property (see Remark 2.8), and for such functions the following holds:
    $F_{\varepsilon}\in W^{j,1}(\Omega),\quad\operatorname{div}(xF_{\varepsilon})\in W^{j,1}(\Omega)\implies F_{\varepsilon}\in W^{j+1,1}(\Omega)\qquad\text{ for all }j\geq 0$
    Indeed, since $\operatorname{div}(xF_{\varepsilon})=dF_{\varepsilon}+x\cdot\nabla F_{\varepsilon}$ and, by hypotheses $x\cdot F_{\varepsilon}\in W^{1,1}(\Omega)$, due to radial symmetry we can write
    $x\cdot\nabla F_{\varepsilon}=x\cdot F_{\varepsilon}^{\prime}(|x|)\frac{x}{|x|}=F_{\varepsilon}^{\prime}(|x|)|x|.$
    Since $0\notin\Omega$, (i.e.$|x|\geq\delta$), it follows that $F^{\prime}(|x|)\in W^{1,1}(\Omega)$. By definition of $\nabla F_{\varepsilon}$ and again the fact that $0\notin\Omega$, we conclude that $\nabla F_{\varepsilon}\in W^{j,1}(\Omega)$, which implies that $F_{\varepsilon}\in W^{j+1,1}(\Omega)$.
@@ 5.2 Moments and cumulants
(always well-defined but possibly +$\infty$+ ∞) and the $\alpha$-moment by
@@
$\mathfrak{M}[f](\xi):=\int_{\mathbb{R}^{d}}e^{x\xi}u(t,x)\,\mathrm{d}x$
whenever $x\mapsto e^{x\xi}u(t,x)$ is integrable. Moments can then also defined via the moment generating function by
@@
If two distributions have the same moments than they have the same cumulants and vice versa. However, moments always depend on the moments of lower order while the cumulants don’t.
@@
Analysing the first term we get
@@
    where $B_{n}$ are the complete (exponential) Bell’s Polynomials defined in (26) and $B_{n,k}$ are called partial Bell’s Polynomials [^1] From a combinatorial point of view $B_{n,k}(x_{1},\dots,x_{n-k+1}$) enumerate the possible ways to partition a set of $n$ elements into $k$ non-empty subsets. Each term of the polynomials express specific way to distribute the $n$ elements in the $k$ subsets. For example, $B_{5,2}=10x_{2}x_{3}+5x_{1}x_{4}$ means that there are $10$ ways to partition a set of $5$ elements into two subsets of cardinality respectively $2$ and $3$ and $5$ ways of doing so in two subsets of cardinality $1$ and $4$
@@
Expression (27) comes directly from the definition of Bell’s olynomials (26). ∎
@@
One can generalise the last result to higher dimension but it appears that the multidimensional version of Bell’s polynomials is not very well known. We refer to Withers and Nadarajah (2010), Schumann (2019) and references therein for further information.
@@
With a direct computation one can obtain a recursive formula for the moments of the solution (and consequently for the equilibrium)
@@
This can be obtained by looking at the time derivative of the $\alpha$ moment, given by:
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}m_{\alpha}(u(t))=\frac{1}{\varepsilon^{2}}(m_{\alpha}(J_{\varepsilon}*u)-m_{\alpha}(u))-m_{\alpha}(\operatorname{div}(xu)).\end{split}$
@@ 5.3 Tails
For the usual (local) equilibrium for the Fokker-Planck equation (FP), is a standard Gaussian. In the nonlocal approximation in this paper, the tails of the equilibrium seem to decay slower, no matter how fast the kernel $J$ is decaying. As a consequence of the Lyapunov condition (Theorem 3.4), for an exponentially decaying kernel $J\in L_{\exp,a^{\prime}}$, the unique probability equilibrium has exponentially decaying tails; that is, for every $a<a^{\prime}\varepsilon^{-1}$
$\int F_{\varepsilon}(x)e^{a\left\langle x\right\rangle}\,\mathrm{d}x<\infty.$
@@
$\int F_{\varepsilon}(x)e^{a\left\langle x\right\rangle^{\gamma}}\,\mathrm{d}x=\infty\quad\text{ for every }\gamma>1.$
@@
Let $\varphi(x)=\exp(a\left\langle x\right\rangle\log\left\langle x\right\rangle)$ with $a$ to be determined later. Following the Lyapunov proof
@@
$\begin{split}\frac{1}{\varepsilon^{2}}&\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\left\langle x+y\right\rangle^{k}-\left\langle x\right\rangle^{k})\,\mathrm{d}y\leq\frac{C_{1}}{2}a^{2}\int_{B_{R}}J(z)\log^{2}\left\langle\xi\right\rangle\varphi(\xi)\,\mathrm{d}z\\ &\leq C_{2}a^{2}\log^{2}\left\langle x\right\rangle\varphi(|x|+\varepsilon R)\int_{B_{R}}J(z)\,\mathrm{d}z=C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\left\langle|x|+\varepsilon R\right\rangle}\\ &\leq C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\varepsilon R}e^{a\left\langle x\right\rangle\log\left\langle x\right\rangle}=\Big{[}C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\varepsilon R}\Big{]}\varphi(x)\end{split}$
@@
$\begin{split}x\cdot\nabla\varphi&=x\cdot a\frac{x}{\left\langle x\right\rangle}\varphi(\log\left\langle x\right\rangle+1)\leq\Big{[}C_{3}a\left\langle x\right\rangle\log\left\langle x\right\rangle\Big{]}\varphi(x)\end{split}$
@@
$\begin{split}L^{*}_{\varepsilon}&\left\langle x\right\rangle^{k}=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y-x\nabla\varphi\\ &\leq\Big{[}C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\varepsilon R}-C_{3}a\left\langle x\right\rangle\log\left\langle x\right\rangle\Big{]}\varphi(x)\leq C-\lambda\varphi\end{split}$ (28)

### v3

@@ 4 Convergence: from nonlocal to local
We leave here a standard result of the decay of the derivatives of the solution of the Fokker-Planck equation.
Lemma 4.5.
Let $u$ be the solution of Fokker-Planck equation with initial data $u_{0}$, $K\geq 0$ and $\alpha$ a multiindex. If $u_{0}$ is such that
$\left|D^{\alpha}u_{0}(x)\right|\leq c\left\langle x\right\rangle^{-K}$
then for every $T^{*}>0$ there exists a constant $C_{T^{*}}$ depending only on $T^{*}$, $K$, $\alpha$ and the dimension $d$ such that
$\left|D^{\alpha}u(x)\right|\leq C_{T^{*}}\left\langle x\right\rangle^{-K}\qquad\text{for all $0\leq t\leq T^{*}$.}$
Sketch of proof.
Since $u(t,x)=e^{dt}v(\frac{1}{2}(e^{2t}-1),xe^{t})$ where $v$ is a solution to the standard heat equation with the same initial data, it is easier to prove the corresponding result for $v$. Since
$v=\Gamma_{t}*u_{0},$
where $\Gamma_{t}$ is the standard heat kernel, passing the derivative to the initial data one easily obtains the result for $v$. Then going back to $u$ one finds a time dependent constant $C_{t}$ such that
$\left|D^{\alpha}u(x)\right|\leq C_{t}\left\langle x\right\rangle^{-K}$
Taking the supremum over $0,T^{*}]$ the proof is complete. ∎
@@
By Lemma 4.5, the solution $v=v(t,x)$ of (FP) is in $\mathcal{C}^{4}$, for every $t\in[0,T^{*}]$ and $|D^{\alpha}v(t,x)|\leq c\left\langle x\right\rangle^{-K}$ uniformly in $0,T^{*}]$, due to our assumptions on the derivatives of $u_{0}$. Hence, $v$ satisfies the hypotheses of Proposition 4.1, uniformly for all $t\in[0,T^{*}]$. We rewrite
$L_{0}v=L_{\varepsilon}v+(L_{0}v-L_{\varepsilon}v):=L_{\varepsilon}v+h(t).$
@@
Proposition 4.7 (Convergence of the equilibrium).
@@
$\left\|S_{t}\right\|\leq Me^{-\gamma t}\qquad\text{ for }\gamma>0,M>1,t>0$ (24)
(where $\left\|\cdot\right\|$ denotes the operator norm) is equivalent to the property that for every $\lambda>-\gamma$, $\lambda$ is in the resolvent set $\rho(L_{\varepsilon})$ of its infinitesimal generator $(L_{\varepsilon},\mathcal{D}(L_{\varepsilon}))$, and
@@
where $R$ denotes the resolvent operator. From Theorem 1.1—in particular from Rmk. 1.2—the condition (24) is satisfied in the Banach space $\mathcal{Y}:=\{f\in L^{1}_{k}:\int_{\mathbb{R}^{d}}f=0\}$. Thus, choosing $n=1$ and $\lambda=0>-\gamma$ in (25), it follows that $L_{\varepsilon}$ has an inverse in $\mathcal{Y}$ with $\left\|L_{\varepsilon}^{-1}\right\|_{\mathcal{Y}}\leq M/\gamma$. This implies that,
$\begin{split}\left\|F_{\varepsilon}-G\right\|_{L^{1}_{k}}&=\left\|F_{\varepsilon}-G\right\|_{\mathcal{Y}}=\left\|L_{\varepsilon}^{-1}L_{\varepsilon}(F_{\varepsilon}-G)\right\|_{\mathcal{Y}}\leq\left\|L_{\varepsilon}^{-1}\right\|_{\mathcal{Y}}\left\|L_{\varepsilon}(F_{\varepsilon}-G)\right\|_{\mathcal{Y}}\\ &\leq\frac{M}{\gamma}\left\|L_{\varepsilon}(F_{\varepsilon}-G)\right\|_{L^{1}_{k}}=\frac{M}{\gamma}\left\|(L_{\varepsilon}-L_{0})G\right\|_{L^{1}_{k}},\end{split}$
since $L_{\varepsilon}F_{\varepsilon}=L_{0}G=0$. Given the regularity and decay properties of the Gaussian, by consistency 4.1 we obtain
@@ 5.1 Regularity
Thanks to the explicit form (11) of the equilibrium in the frequency domain, one can study the regularity by standard results on the Fourier transform.
Recall that the Fourier transform of the equilibrium is
with $\zeta_{\varepsilon}(\xi)=\hat{J}_{\varepsilon}(\xi)-1$.
@@
Let us assume $\zeta_{\varepsilon}$ radially symmetric. If not, define $\underline{\zeta}_{\varepsilon}$ and $\overline{\zeta}_{\varepsilon}$ by
@@
Since $\zeta_{\varepsilon}(y)=-\varepsilon^{2}y^{2}+o(y^{2})$, the singularity at zero is canceled and, moreover, close to $\xi=0$ we have
@@
    which implies that |^F_ε(ξ)|≤Cexp(-1-δε2log|ξ|)=C|ξ|1-δε2C|ξ|k+(1ε2-k)/2. For $|\xi|$ large, the previous implies condition (26) with $\eta<(\frac{1}{\varepsilon^{2}}-k)/2$ and by standard argument on Fourier transform $F_{\varepsilon}\in\mathcal{C}^{k-d}$ for all $k<\frac{1}{\varepsilon^{2}}$.
@@
    proceeding in a similar way as before. Notice that since $\xi\mapsto\hat{J}(\xi)$ is continuous and decays to zero as $|\xi|\to+\infty$, there exists $y_{\delta}$ such that we can bound $\hat{J}$ from below by ^J(x)≥- δ:=ε22(1ε2-k)   for all y≥y_δ. As a consequence we see that
@@
We just proved that the regularity keeps improving when $\varepsilon\to 0$ although the result is not optimal. A precise characterization remains, to our knowledge, an open question. Moreover, we suspect that under quite general conditions, the only point at which $F_{\varepsilon}$ is not $\mathcal{C}^{\infty}$ is the origin. We are able to show this in dimension $1$, and for radially symmetric $J$:
@@
    If $d\neq 1$, the fact that $\operatorname{div}(xF_{\varepsilon})$ is integrable in $\Omega$ does not assure that $\nabla F\in L^{1}(\Omega)$, even when we are far from 0. Therefore, we need an additional hypothesis. If $J$ is radially symmetric, $F_{\varepsilon}$ inherits this property (see Remark 2.8), and for such functions the following holds: F_ε∈W^j,1(Ω), div(xF_ε)∈W^j,1(Ω) ⟹F_ε∈W^j+1,1(Ω)   for all j≥0 Indeed, since $\operatorname{div}(xF_{\varepsilon})=dF_{\varepsilon}+x\cdot\nabla F_{\varepsilon}$ and, by hypotheses $x\cdot F_{\varepsilon}\in W^{1,1}(\Omega)$, due to radial symmetry we can write x⋅∇F_ε= x⋅F_ε’(|x|)x|x|=F_ε’(|x|)|x|.
[…]
@@ 5.2 Moments and cumulants
(always well-defined but possibly $\infty$) and the $\alpha$-moment by
@@
$\mathfrak{M}[f](\xi):=\int_{\mathbb{R}^{d}}e^{x\xi}f(t,x)\,\mathrm{d}x$
whenever $x\mapsto e^{x\xi}u(t,x)$ is integrable. Moments can then also be defined via the moment generating function by
@@
If two distributions have the same moments then they have the same cumulants and vice versa. However, moments always depend on the moments of lower order while the cumulants don’t.
@@
Analyzing the first term we get
@@
    where $B_{n}$ are the complete (exponential) Bell’s Polynomials defined in (28) and $B_{n,k}$ are called partial Bell’s Polynomials [^1] From a combinatorial point of view, $B_{n,k}(x_{1},\dots,x_{n-k+1}$) enumerates the possible ways to partition a set of $n$ elements into $k$ non-empty subsets. Each term of the polynomials represents specific way of distributing the $n$ elements in the $k$ subsets. For example, $B_{5,2}=10x_{2}x_{3}+5x_{1}x_{4}$ means that there are $10$ ways to partition a set of $5$ elements into two subsets of cardinality respectively $2$ and $3$ and $5$ ways to do so in two subsets of cardinality $1$ and $4$. .
@@
Expression (29) comes directly from the definition of Bell’s polynomials (28). ∎
@@
One can generalize the last result to higher dimension but it appears that the multidimensional version of Bell’s polynomials is not very well known. We refer to Withers and Nadarajah (2010), Schumann (2019) and references therein for further information.
@@
By computing explicitly the time derivative of the $\alpha$ moment, i.e.
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}m_{\alpha}(u(t))=\frac{1}{\varepsilon^{2}}(m_{\alpha}(J_{\varepsilon}*u)-m_{\alpha}(u))-m_{\alpha}(\operatorname{div}(xu)),\end{split}$
one can obtain the following recursive formula for the moments of the solution (and consequently for the equilibrium):
@@ 5.3 Tails
For the usual (local) equilibrium for the Fokker-Planck equation (FP), is a standard Gaussian. In the nonlocal approximation in this paper, the tails of the equilibrium seem to decay slower, no matter how fast the kernel $J$ is decaying. As a consequence of Harris’s theorem and our Lyapunov condition estimates (Theorem 1.3 (ii)), we know for an exponentially decaying kernel $J\in L_{\exp,a^{\prime}}$, the unique probability equilibrium has exponentially decaying tails; that is, for every $a<a^{\prime}\varepsilon^{-1}$
$\int_{\mathbb{R}^{d}}F_{\varepsilon}(x)e^{a\left\langle x\right\rangle}\,\mathrm{d}x<\infty.$
@@
$\int_{\mathbb{R}^{d}}F_{\varepsilon}(x)e^{a\left\langle x\right\rangle^{\gamma}}\,\mathrm{d}x=\infty\quad\text{ for every }\gamma>1.$
@@
Let $\varphi(x)=\exp(a\left\langle x\right\rangle\log\left\langle x\right\rangle)$ with $a$ to be determined later. In the same spirit of the proofs of theorems 3.2 and 3.4, one can compute
@@
$\begin{split}\frac{1}{\varepsilon^{2}}&\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y\leq\frac{C_{1}}{2}a^{2}\int_{B_{R}}J(z)\log^{2}\left\langle\xi\right\rangle\varphi(\xi)\,\mathrm{d}z\\ &\leq C_{2}a^{2}\log^{2}\left\langle x\right\rangle\varphi(|x|+\varepsilon R)\int_{B_{R}}J(z)\,\mathrm{d}z=C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\left\langle|x|+\varepsilon R\right\rangle}\\ &\leq C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\varepsilon R}e^{a\left\langle x\right\rangle\log\left\langle x\right\rangle}=\Big[C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\varepsilon R}\Big]\varphi(x)\end{split}$
@@
$\begin{split}x\cdot\nabla\varphi&=x\cdot a\frac{x}{\left\langle x\right\rangle}(\log\left\langle x\right\rangle+1)\varphi(x)\leq\Big[C_{3}a\left\langle x\right\rangle\log\left\langle x\right\rangle\Big]\varphi(x)\end{split}$
@@
$\begin{split}L^{*}_{\varepsilon}&\varphi(x)=\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\varphi(x+y)-\varphi(x))\,\mathrm{d}y-x\cdot\nabla\varphi(x)\\ &\leq\Big[C_{2}a^{2}\log^{2}\left\langle x\right\rangle\left\langle|x|+\varepsilon R\right\rangle^{a\varepsilon R}-C_{3}a\left\langle x\right\rangle\log\left\langle x\right\rangle\Big]\varphi(x)\leq C-\lambda\varphi\end{split}$ (30)